Consider the following information:

  

Rate of Return if State Occurs
  State of Economy Probability of State of Economy Stock A Stock B
  Recession 0.10 0.05 -0.17
  Normal 0.50 0.07 0.11
  Boom 0.40 0.14 0.33

  

Required:
(a)

Calculate the expected return for Stock A. (Do not round your intermediate calculations.)

(Click to select)9.60%8.19%11.33%10.37%8.73%

  

(b)

Calculate the expected return for Stock B. (Do not round your intermediate calculations.)

(Click to select)17.00%9.00%18.85%16.15%17.68%

  

(c)

Calculate the standard deviation for Stock A. (Do not round your intermediate calculations.)

(Click to select)3.64%2.57%3.82%3.46%3.78%

  

(d)

Calculate the standard deviation for Stock B. (Do not round your intermediate calculations.)

(Click to select)15.36%10.86%17.13%14.59%15.98%
Answer #1

a. Calculation of Expected Return for Stock A: = (0.10 x 0.05)+(0.50x0.07)+(0.40 x 0.14) = 0.005 + 0.035 + 0.056 = 9.60% a. C